Help I Given the accounting equation, answer each of the following questions. pany are $126,500 and its stockholders equity is $222,500. What is the amount of Holland Companys total assets? Question
Answer #1

a) Total assets = Total liabilities+Total Stockholder's equity

= 126500+222500

Total assets = $349000

b) Total liabilities = Total assets-Total stockholder's equity

= 187500-93000

Total liabilities = $94500

c) Owner's equity = Total assets-Total liabilities

= 884500-(884500/2)

Owner's equity = $442250